# User Simulation for Evaluating Information Access Systems

### v1

@@ 7.5.2 Conversational Information Access
A series information seeking benchmark datasets are developed within the TREC 2019–2022 Conversational Assistance track (TREC CAsT) [Dalton:2020:SIGIR]. [^18] https://www.treccast.ai/ Resources for clarifying questions include Qulac [Aliannejadi:2019:SIGIR] and its extension ClariQ [Aliannejadi:2020:arXiv]; both comprise (topic, facet, clarifying question, answer) tuples for single-turn interaction. A multi-turn dataset is created and released in [Sekulic:2022:WSDM]. There are several conversational recommendation datasets available for the movies domain, e.g., the Movie Dialogue datasets [Dodge:2016:ICLR], [^19] https://research.facebook.com/downloads/babi/ ReDial [Li:2018:NIPS], [^20] https://redialdata.github.io/website/ and INSPIRED [Hayati:2020:EMNLP]. [^21] https://github.com/sweetpeach/Inspired DuRecDial [Liu:2021:EMNLP] [^22] https://github.com/liuzeming01/DuRecDial covers additional domains as well, including music, food, and restaurants. MG-ShopDial is multi-goal dataset for conversational information access that includes conversational search, recommendation, and QA at the same time, for the e-commerce domain [Bernard:2023:SIGIR]. [^23] https://github.com/iai-group/mg-shopdial
7.6 Validating Simulators
Validation concerns the evaluation of the user simulator itself. Pietquin:2013:KER identify desired features for a good user simulation metric. In doing so, they emphasize that “it is necessary to provide a clear idea of the purpose of a user simulation” [Pietquin:2013:KER]; there are different requirements for simulators that are used for training RL-based dialogue managers, expanding existing datasets, or facilitating automatic evaluation of systems. Our interest is in the latter category. We discuss the evaluation of user simulators for conversational information access tasks on three different levels: (1) individual utterances, (2) individual dialogues, and (3) a collection of generated dialogues.
For utterance-level evaluation, it is common to have human raters evaluate the generated responses along different dimensions, such as naturalness, usefulness, and grammar [Sekulic:2022:WSDM, Zhang:2022b:SIGIR, Owoicho:2023:SIGIR].
For evaluating entire dialogues, Zhang:2020:KDD propose a side-by-side human evaluation protocol, where assessors are given transcripts of two conversations, in random order, and they are told that one of the dialogues was conducted by a human and the other one by a “bot.” They have to guess which of the two is the human one. This follows the idea of adversarial evaluation [Li:2017:EMNLP], inspired by the Turing test, where the more successful a simulator in “fooling” the human evaluator, the more realistic it is.
A collection of simulated dialogues may be compared against a human-machine dialogue corpus “to evaluate how well the simulated dialogues match the statistical distribution of the real dialogue corpus” [Schatzmann:2006:KER]. Various corpus-level statistics have been proposed in the literature [Schatzmann:2005:SIGDIAL, Schatzmann:2006:KER, Pietquin:2013:KER] that can be used as automatic measures. These may be divided into three groups: (1) high-level dialogue features, such as average dialogue length (number of turns) and ratio of user vs. system actions; (2) dialogue style, which includes the frequency or distribution of dialogue acts, user cooperativeness (proportion of slot values provided when requested); and (3) dialogue efficiency, including success (or task completion) rate, reward, and completion time.
The ultimate measure of success of a user simulator is how well it can predict the performance of a system with real users. This cannot be carried out without the involvement of real users. Thus, the same system(s) need to be evaluated using both human and simulated users. The obtained measures may be compared in absolute terms. Often, however, we are interested in a relative comparison between systems, in which a case it is sufficient to consider the relative ordering of systems [Zhang:2020:KDD].
7.7 Summary and Future Challenges
Research on conversational information access systems is still in a relatively early phase. Much of the research has focused on individual components and highly goal-specific dialogue flows. There are user simulators available to support the automatic evaluation of these scenarios, but it is also clear that there is much headroom for improving the realisticity and robustness of these simulators.
One of the biggest challenges is that the evaluation of conversational systems itself is—still—an open question. Even though a number of different evaluation measures have been proposed [Deriu:2021:AIR, Liu:2021:TOIS], these fail to capture what makes a conversation ultimately successful. As Zamani:2023:arXiv explain, “the same conversation may be considered of high or of low quality depending on context: For example, if a user is in a rush or not, or if the user requires high confidence in the conclusion or not.” Shorter time to task completion is not necessarily sufficient; a prolonged but engaging interactions may lead to a more satisfying overall experience. Also, users might trade off long- and short-term utility [Radlinski:2017:CHIIR].
In addition to those, facilitating multi-modal interactions would be especially pertinent in a conversational setting. While past research has primarily focused on uni-modal (i.e., natural language) interactions, other modalities as input (e.g., speech, pointing/clicking, body gestures) and output (e.g., speech and multimedia elements) have also begun to receive increasing attention [Liao:2021:SIGIR, Kostric:2022:RecSys, Deldjoo:2021:SIGIR, Hauptmann:2020:MM]. Modeling the action space around these and selecting the appropriate modality for a given user action/context are interesting challenges to be addressed in future simulators.
Furthermore, progress on conversational information access systems that support and seamlessly integrate multiple users goals critically depends on appropriate evaluation methodology and resources [Balog:2021:DESIRES]. User simulation techniques play a pivotal role in fostering research and development in this area.
Chapter 8 Conclusion and Future Challenges
@@ 7.7 Summary and Future Challenges
Accurate evaluation of information access systems—such as search engines, recommender systems, and conversational assistants—is critical both for confident real-world deployment and for making progress in scientific research. However, the task of effectively evaluating these systems has long remained a challenging endeavor. At least one major reason for the difficulty lies in the fact that the utility of such a system generally has to be assessed by its users, and different users (or even the same user in a different use context) often perceive the utility differently. This also reflects the general challenge of evaluating any interactive intelligent system whose utility must be assessed by its user in a context-specific way.
@@
As one promising way to address this challenge, user simulation has been proposed and studied in the context of information access systems, with much progress made in the last two decades. We started the book with a broad overview of user simulation and its many applications with its use for evaluation as our focus. We showed that current evaluation methods and measures can generally be viewed as naive user simulators. Thus, user simulation naturally serves as a unified framework for developing better evaluation methods in the future. Indeed, user simulation may also be the only way to support reproducible experiments on evaluating any interactive intelligent system whose utility must be assessed by its users empirically. We then systematically reviewed the frameworks, models, and techniques for constructing user simulators by using an MDP framework to tie multiple lines of scattered research work together in a unified formal framework. We now conclude this survey by highlighting open issues and providing several potential research directions. We end with some recommendations of actions for the broader research communities.
@@ 8.1 Embracing Simulation-based Evaluation
There has been significant progress in the development of user simulation techniques, and we have seen numerous instances of successful application of simulation-based approaches for evaluating search engines, recommender systems, and conversational assistants. The benefits of using user simulation for system evaluation are numerous. Simulation allows researchers and developers to test their systems under various scenarios and conditions, which may be difficult or impossible to achieve in real-world testing. Additionally, user simulation can help identify potential flaws or weaknesses in a system before it is deployed, leading to more robust and effective systems.
@@
Despite its clear benefits, simulation-based evaluation has not been widely adopted in the information retrieval and recommender systems communities. This may be due to several factors, including the complexity of creating realistic simulations, the lack of consensus on simulation-based evaluation methodology, open questions regarding the validity of simulations, and the resources required to develop and run simulations. We argue that it is time to recognize the value and benefits of user simulation and that it should be considered as a valuable addition to the evaluation toolbox of researchers and developers, alongside with the methodologies of test collection-based evaluation, user studies, and online evaluation. Indeed, by viewing an existing evaluation measure using the user simulation evaluation framework, we can often obtain a more meaningful interpretation of the evaluation measure and assess any unrealistic assumption about a user implicitly made by the evaluation measure. [^1] One specific example is the natural justification in the user simulation framework for the use of variable $k$ when using NDCG@k as a measure to aggregate over a set of queries [Karmaker:2020:CIKM].
@@
To promote user simulation for evaluation, as a first step, we should leverage the many test collections already available, notably those created at TREC and CLEF, to turn those into user simulators that are made publicly available to the community. Even though present-day user simulators are far from perfect or accurate, compared with the current evaluation methods used (where unrealistic or extremely naive user behaviour has been assumed), using user simulation for evaluation is still desirable, let alone the opportunity to evaluate interactive systems that we currently cannot easily do without relying on user simulation. [^2] Note that despite the incompleteness of relevance judgments, relative comparison of systems using such imperfect relevance judgments is still meaningful, so in this sense, the community is already accepting this kind of “imperfect evaluation.”
@@
User simulators can be released as toolkits, such as SimIIR [Maxwell:2016:SIGIR, Zerhoudi:2022:CIKM] or UserSimCRS [Afzali:2023:WSDM], or a web-based API service as has been done in living lab-style evaluation [Hopfgartner:2019:bookchapter, Jagerman:2018:JDIQ]. These relevant existing efforts have already demonstrated the feasibility of using more interactive evaluation approaches. To promote the use of user simulation for evaluation, it is also necessary to organize evaluation activities regularly (e.g., by establishing a user simulation track at TREC or CLEF) to facilitate research and sharing of knowledge.

### v2

@@ 8.4.3 Conversational Assistants
A series of information seeking benchmark datasets are developed within the TREC 2019–2022 Conversational Assistance track (TREC CAsT) [Dalton:2020:SIGIR]. [^17] https://www.treccast.ai/ Resources for clarifying questions include Qulac [Aliannejadi:2019:SIGIR] and its extension ClariQ [Aliannejadi:2020:arXiv]; both comprise (topic, facet, clarifying question, answer) tuples for single-turn interaction. A multi-turn dataset is created and released in [Sekulic:2022:WSDM]. There are several conversational recommendation datasets available for the movies domain, e.g., the Movie Dialogue datasets [Dodge:2016:ICLR], [^18] https://research.facebook.com/downloads/babi/ ReDial [Li:2018:NIPS], [^19] https://redialdata.github.io/website/ and INSPIRED [Hayati:2020:EMNLP]. [^20] https://github.com/sweetpeach/Inspired DuRecDial [Liu:2021:EMNLP] [^21] https://github.com/liuzeming01/DuRecDial covers additional domains as well, including music, food, and restaurants. MG-ShopDial is multi-goal dataset for conversational information access that includes conversational search, recommendation, and QA at the same time, for the e-commerce domain [Bernard:2023:SIGIR]. [^22] https://github.com/iai-group/mg-shopdial
Chapter 9 A Broader Perspective on User Simulation
In this book so far, we have primarily focused on using user simulation as an evaluation tool for information access systems, such as search engines, recommender systems, and conversational assistants. However, the applications of simulation extend to assistive AI systems in general and far beyond evaluation. Simulation techniques can be employed for data augmentation, generating synthetic user interactions to improve the training of machine learning models, or user behaviour modeling, and creating realistic simulations of how users interact with a system in general (Section 9.1). As a research topic, user simulation is inherently interdisciplinary, intersecting with diverse fields both within and beyond computer science. For example, it draws upon concepts from psychology, economics, and human-computer interaction to create accurate and representative models of user behaviour (Section 9.2). The recent success of large language models (LLMs) has led to their widespread use in simulation tasks across different domains and application scenarios, from generating realistic dialogue for chatbots to simulating customer interactions for business training (Section 9.3). Finally, we take an even broader perspective of user simulation to discuss its connection with Artificial General Intelligence (AGI), the long-term goal of creating AI systems with human-like intelligence (Section 9.4).
9.1 Broad Applications of User Simulation
Although our focus in the book is on the use of simulation for evaluation, user simulation has numerous other applications, including particularly user behaviour modeling and analysis, and data augmentation, which we briefly discuss in this section. Different applications may require or appreciate different types of simulation techniques, but there is clear synergy of those different lines of exploration. Each application area will likely lead to the development of distinct simulators. Multiple lines of research can mutually enhance and inform each other. Eventually, these diverse lines of research may even converge as we progress towards AGI.
9.1.1 Data Augmentation
There has been much work on using user simulation for data augmentation. The generated synthetic data can then be used for training machine learning algorithms. While interpretability is critical when we use user simulation for evaluation, it is not so important when using it for data augmentation as whether the generated data would be useful might not depend so much on the realism of a simulator, but rather whether the distribution of new data is correlated with the distribution of true user data. Such a relaxed requirement makes it easier to make progress in user simulation for data augmentation especially because of the emergence of deep learning techniques that enable end-to-end training of an algorithm for virtually any prediction problem (predicting user behaviour would enable data augmentation). The availability of LLMs further makes it feasible to simulate users to generate realistic text content, which opens up even more opportunities for using user simulation for data augmentation.
When we use a user simulator to evaluate an interactive system, the evaluation process would also naturally generate additional interaction data of the simulated users. Such data can generally be leveraged for data augmentation. When a user simulator is interpretable, the user simulator can be varied via changing meaningful user variables to simulate potentially many more users in a counterfactual way (e.g., we could make a simulated user increasingly more patient) and thus is, in general, more powerful than a non-interpretable simulator.
While there are many ways to generate synthetic data for data augmentation, a major advantage of data generated by a user simulator is that the data reflect a user’s interaction with a system, which are especially useful for training an interactive intelligent system. Such data, along with evaluation measures defined on user interactions, can provide reward information for an intelligent interactive system to use for optimizing its interaction policy, thus enabling end-to-end training of a whole interactive agent.
However, it is also worth pointing out that although simulating a whole user is desirable, for data augmentation, even simulating a single user behaviour (e.g., clicking in search) would still be useful for obtaining additional training data. Focusing on specific actions would also simplify making progress in data augmentation. Since data augmentation benefits various machine learning models, using user simulation for this purpose further expands the scope of research on user simulation into potentially many other areas of AI.
9.1.2 User Behaviour Analysis
Analysis of user behaviour is important for many reasons. For example, it enables understanding users in general, which further enables understanding our society and and answering many social science research questions. Also, formal user models to capture user behaviour can be used by an intelligent agent to personalize its service to individual users.
User simulation can be regarded as taking a computational approach to studying users’ behaviour. Specifically, once we build a user simulator, the simulator can be regarded as a formal (mathematical) model of a user, where all the hypotheses about the user’s behaviour are explicitly articulated in some form, thus enabling us to test those hypotheses by comparing the behaviour of a simulated user with that of a real user. Such a comparison can directly facilitate validation of a user simulator and provide insights about how to further improve a simulator. For example, we may discover the simulator is unable to simulate a particular behaviour of a user well, and thus consider this weakness when using the user simulator for evaluating systems. Such insights can also guide further research on improving user simulation.
Analysis of user behaviour has often been done by doing controlled user studies. Such work often reveals interesting observations or insights about behaviours of real users that can be leveraged directly for building user simulators. In this sense, such user studies can be viewed part of the general research agenda of building realistic user simulators, only that they may focus on simulating one aspect of user behaviour in a particular user study. From another perspective, user simulation enables testing any hypothesis that might be formulated based on any real user study.
Analysis of user behaviour has also been done by analyzing log data generated by real users. Such work can be regarded as directly contributing to research on user simulation, especially when we can also use such log data for instantiating or seeding a user simulation (see more discussion of this issue in Section 8.1). Indeed, it is natural to view all work on analyzing log data of users as part of study of user simulation, thus the scope of research in user simulation generally covers analysis of all kinds of user interaction logs in any system, which does not necessarily have to be an AI system.
An especially interesting direction of using user simulation for user modeling is to conduct simulation experiments to model a community of users interacting with an interactive system over a long time period as has been done in, e.g., [Hazrati:2024:UMUAI, Yao:2023:ICML]. Such simulation experiments would enable counterfactual analysis of impact of any system to be deployed as well as understanding user behaviour and preferences in a simulated application environment such as recommender systems [Zhang:2020:ISR].
9.2 User Simulation as an Interdisciplinary Research Field
The broader discussion of user simulation in the previous section also suggests that the research of user simulation is a highly interdisciplinary and touches many different topic areas, including, e.g., information science, information access, machine learning, natural language processing, knowledge representation, human-computer interaction, and psychology. It is thus desirable to integrate research in all these different areas in the future. In the rest of this section, we will discuss some specific connections between user simulation with multiple other areas.
9.2.1 User Simulation and Intelligent Agents
A sophisticated user simulator can be regarded as an intelligent agent that operates in an environment of a system by interacting with the system. The MDP framework that we discussed earlier provides a foundation for formally defining the behaviour of such an agent. As such, research on user simulation requires studying how to build an intelligent agent in general and many techniques used for building an intelligent agent are also useful for building a user simulator. We thus anticipate increasingly more such techniques to be used for user simulation. From another perspective, research on user simulation enables intelligent agents to interact with simulated users for training their interaction policy, thus enabling improvement of interactive intelligent agents.
Considering human-AI collaboration, we can expect sophisticated user simulators to not only simulate human-like behaviour but also simulate/model human’s ability of problem solving since in the future, an intelligent agent may (proacively) ask a human user for help, in which case, the user simulator would need to simulate what kind of help the user may potentially provide to an AI system. Since it is also desirable to make an intelligent (task) agent human-like, building user simulation agents and building intelligent task agents may be viewed as approaching the same goal of building an intelligent human-like agent from different angles, both moving toward artificial general intelligence as we will further discuss in Section 9.4
9.2.2 User Simulation and Machine Learning
User simulation provides many interesting opportunities for applying a broad variety of ML techniques, including unsupervised learning, supervised learning, imitation learning, and reinforcement learning (RL). Particularly relevant is inverse reinforcement learning [Arora:2021:AI], as it is especially designed to recover the reward function and the policy just from the analysis of the observed behaviour. This is unlike RL where access to the reward function is assumed, which is often difficult to define. Generative probabilistic models may be applied to learn patterns of user actions in an unsupervised way. The observed user actions can be used as labels for use with supervised learning or imitation learning to build a user simulator. Reinforcement learning is necessary for user simulation since it enables a user simulator to optimize sequential decision making (policy) by learning from its interactions with a system. User simulation can also be used to generate synthetic user (interaction) data which can be useful for training machine learning algorithms, especially complex algorithms, such as reinforcement learning, that often requires either massive amounts of user interaction data or an interactive environment for training. User simulation techniques would enable the training of such algorithms that are currently infeasible.
All such knowledge also needs to be updated to reflect learning by a search user during the search process [Vakkari:2016:JIS]. Therefore, it is necessary to study the cognitive processes of searchers. This makes the study of user simulation closely related to intelligent tutoring systems, where simulated learners have also been used for evaluating intelligent education systems [Dorcca:2015:AIED]. In general, how to represent the knowledge that a user possesses, how to model the reasoning a user makes over the knowledge, and how to capture a user’s knowledge gain during a search process are all important yet challenging open questions that may require deeper understanding of a fundamental question in AI—how knowledge should be represented formally.
9.2.4 User Simulation and HCI and Psychology
Psychology and HCI research play a vital role in establishing the theoretical foundations for simulating users, making them not only closely linked to user simulation but also essential for constructing realistic simulators. Looking ahead, we anticipate a growing integration of knowledge and theories from psychology and HCI into the development of user simulators. Leveraging any insights gained about users has the potential to enhance the accuracy and fidelity of user simulation. Starting points might include Fu:2007:HCI’s work on SNIF-ACT, well-motivated and explicit computational cognitive model that explains navigation behaviour on the Web [Fu:2007:HCI]. Another excellent example of a user simulator modeling user surfing behaviour can be found in [Chi:2003:CHI].
9.2.5 User Simulation and Specific User Tasks
Domain-specific or system-specific user simulation would require collaborative research between user simulation researchers and domain experts. The incorporation of domain-specific knowledge and assumptions would be necessary. Since information access systems are widely used by many people in all kinds of application domains, there will be widespread opportunities for collaboration with domain experts, leading to many specialized user simulators developed for specific domains. For example, in the case of simulating searches in the medical domain, we likely would need to incorporate specific knowledge about medical search tasks, whereas in e-commmerce search, it would be necessary to incorporate other kinds of domain knowledge, depending on the specific type of products to be purchased by the user. Simulation of learners in education is another example, which has attracted much attention recently [Kaser:2023:AIED]. Since learners often rely on an information access system to learn, simulation of learners is very closely related to simulation of users of information access systems, and similar issues such as validity of a simulator have also been studied and discussed in the education context [Kaser:2023:AIED].
9.2.6 User Simulation and Software Systems
Finally, in contrast to the static test collections currently used for evaluation, user simulators need to function as operational software platforms to facilitate evaluation. Consequently, an interesting relationship emerges between user simulation and software systems, as the challenge lies in constructing “high-performance” user simulators capable of simultaneously interacting with multiple search systems (to enable their evaluation). Given that a sophisticated user simulator will likely have to use advanced algorithms and data structures (e.g., for representation of knowledge) and support large-scale experiments (that possibly explore various configurations), there will be significant new challenges in building efficient user simulators.
9.3 Large Language Models and User Simulation
LLMs have shown great promise in many AI tasks. With multimodality, the most sophisticated foundation LLMs are now able to simulate many aspects of human intelligence, notably natural language understanding and generation, computer vision, and software coding. While those LLMs are still lacking in their capacity of reasoning and planning and they are also unable to regulate their hallucination behaviours, it is reasonable to assume that they will become increasingly capable and human-like in the future. As such, they provide a good foundation for building both intelligent task agents and user simulation agents. In this section, we briefly discuss some recent trends in using LLMs for user simulation and suggest some promising future directions.
First, since LLMs are able to generate fluent realistic natural language text, they directly enable simulation of all kinds of user actions where the user would produce text, including, e.g., generation of queries for search or the generation of natural language conversational utterances. While simulation of keyword queries has so far been mostly based on simple language models (e.g., n-gram language models) as we discussed in Section 6.2, LLMs have recently been exploited to simulate query generation [Abolghasemi:2023:CIKM, Kiesel:2024:ECIR]. In the case of conversational systems, traditional simulation methods were limited in their capability of generating natural language conversations that a human could produce. LLMs overcame this limitation and can be leveraged to simulate human conversations in various contexts, including, e.g., conversational AI [Meyer:2022:CUI], task-oriented dialogue [Hu:2023:CIKM, Davidson:2023:arXiv], conversational search [Sekulic:2022:WSDM, Owoicho:2023:SIGIR], conversational recommendation [Wang:2023:EMNLP, Friedman:2023:arXiv], and conversational QA [Abbasiantaeb:2024:WSDM].
In addition, LLMs have also been used for simulation in recommender systems [Zhang:2023:arXiv, Wang:2024:arXiv] and for generating relevance label as a way to replace humans in evaluating IR systems [Faggioli:2023:ICTIR]. The latter may be seen as simulating human assessments of utility of a document in search results (using a black box model).
As the uses of LLMs for various information access tasks become more widespread, it is also crucial to acknowledge their limitations. As discussed in Chapter 7, LLM-based responses can be uncontrollable, leading to unrealistic or incoherent behaviours, and a lack of the natural variation found in human interactions [Terragni:2023:arXiv, Davidson:2023:arXiv, Wang:2024:WWW, Yoon:2024:arXiv]. Moreover, LLMs often exhibit greater knowledge than average users and generate overly “perfect” responses, potentially leading to the simulation of unrealistic “superusers” [Wang:2023:EMNLP, Wang:2024:WWW]. While prompting can guide LLM behaviour, ensuring strict adherence to instructions remains a challenge [Kiesel:2024:ECIR, Wang:2024:WWW]. The reproducibility of LLM simulations is also questionable, as model updates may alter behaviour.
LLMs offer great potential for simulating diverse user actions, opening up exciting avenues for future research. For example, LLMs have already been used to build generative agents, i.e., computational software agents that simulate believable human behaviour, enabling various simulations of social environments (e.g., simulating a small town of twenty-five generative agents using natural language [Park:2023:UIST]). For comprehensive surveys of work combining agent-based modeling and LLMs across various fields (social science, natural science, and engineering) and domains (e.g., cyber, physical, social environment) we refer readers to [Gao:2023:arXiv] and [Wang:2024:FCS].
9.4 User Simulation as a Step toward AGI
The overarching goal of developing a realistic user simulator is, in many ways, aligned with the general goal of developing intelligent agents with human-like intelligence, i.e., the goal of achieving Artificial General Intelligence (AGI). A sophisticated user simulator, behaving like a human user, would respond to system interactions and choose appropriate actions to complete a task. As an intelligent agent, such a simulator can be effectively modeled using the MDP framework discussed in this book. Similarly, an intelligent agent designed to support a user in completing a task would also respond to user actions and select appropriate actions (i.e., choose appropriate system services to serve users), and can be generally modeled with an MDP framework as well. Thus, both user simulation agents and intelligent interactive task agents share a common modeling approach, allowing for optimization through reinforcement learning algorithms. The difference lies in the definition of states and actions as well as the reward functions. For user simulation, the primary goal (reward) is to exhibit behaviour similar to real users, while for task agents, the reward is often tied to providing user-friendly service to facilitate task completion. Additionally, the actions differ, with user simulation focusing on user actions and intelligent agents on system responses.
Due to this similarity, the technical challenges in building intelligent user simulators may mirror those faced in developing intelligent task agents. We therefore anticipate broader connections between research in user simulation and many other areas of AI.
We may also view intelligent user simulation agents and intelligent task agents as approaching the same (eventual) goal of AGI from the two ends of a spectrum with variable trade-off between “human-like” and “task support.” At one end of the spectrum, we have intelligent task agents that may emphasize only task support without being human like; at the other end, we have intelligent user simulators that emphasize only human-like without providing task support. Future research can be expected to converge to the middle of the spectrum as those task agents become increasingly human like, while the user simulators become increasingly capable of supporting tasks. The emergence of LLMs may accelerate the integration and synergy discussed above since the foundation LLMs may very well serve as the foundation for building both kinds of agents (eventually).
AI systems can be broadly categorized into two types: autonomous AI, where machines aim to perform tasks normally performed by humans (e.g., self-driving cars), and assistive AI, where the goal is not to replace humans, but rather aid them to fulfill their tasks more efficiently and effectively, often by augmenting human intelligence. A main challenge unique to assistive AI systems is how to collaborate with a human user effectively so as to maximize the combined intelligence of humans and machines. Effective human-AI collaboration would require user simulation agents to be tightly integrated with task agents so that the task agent would be able to leverage a user simulation agent to obtain reward and guide its optimization of interaction policy. Similarly the user simulation agent would also need to adapt to model how new AI systems would behave over time. Thus the dependency between research on intelligent task agents and that on user simulation is inherent and might not disappear until we achieve AGI.
Chapter 10 Conclusion and Future Challenges
@@ 9.4 User Simulation as a Step toward AGI
Accurate evaluation of information access systems—such as search engines, recommender systems, and conversational assistants—is critical both for confident real-world deployment and for making progress in scientific research. However, the task of effectively evaluating these systems has long remained a challenging endeavor. At least one major reason for the difficulty lies in the fact that the utility of such a system generally has to be assessed by its users, and different users (or even the same user in a different use context) often perceive the utility and behave differently. This also reflects the general challenge of evaluating any interactive intelligent system whose utility must be assessed by its user in a context-specific way.
@@
As one promising way to address this challenge, user simulation has been proposed and studied in the context of information access systems, with much progress made in the last two decades. We started the book with a broad overview of user simulation and its many applications with its use for evaluation as our focus. We showed that current evaluation methods and measures can generally be viewed as naive user simulators. Thus, user simulation naturally serves as a unified framework for developing better evaluation methods in the future. Indeed, it is our belief that user simulation may be the only way to support reproducible experiments on evaluating any interactive intelligent system whose utility must be assessed by its users empirically. We then systematically reviewed the frameworks, models, and techniques for constructing user simulators by using an MDP framework to tie multiple lines of scattered research work together in a unified formal framework. We now conclude this survey by highlighting open issues and providing several potential research directions. We end with some recommendations of actions for the broader research communities.
@@ 10.1 Embracing Simulation-Based Evaluation
There has been significant progress in the development of user simulation techniques, and we have seen numerous instances of successful application of simulation-based approaches for evaluating search engines and conversational assistants (though the use of simulation for evaluating recommender systems appears to be less common). The benefits of using user simulation for system evaluation are numerous. Simulation allows researchers and developers to test their systems under various scenarios and conditions, which may be difficult or impossible to achieve in a real-world setting. Additionally, user simulation can help identify potential flaws or weaknesses in a system before it is deployed, leading to more robust and effective systems [Bernard:2024:CUI].
@@
Despite its clear benefits, simulation-based evaluation has not been widely adopted in the information retrieval and recommender systems communities. This may be due to several factors, including the complexity of creating realistic simulations, the lack of consensus on simulation-based evaluation methodology, open questions regarding the validity of simulations, and the resources required to develop and run simulations. Naturally, in the future, we need to do more research to address these challenges. Meanwhile, we also argue that it is time to recognize the value and benefits of user simulation and that it should be considered as a valuable addition to the evaluation toolbox of researchers and developers, alongside with the methodologies of test collection-based evaluation, user studies, and online evaluation. Indeed, by viewing an existing evaluation measure using the user simulation evaluation framework, we can often obtain a more meaningful interpretation of the evaluation measure and assess any unrealistic assumption about a user implicitly made by the evaluation measure. [^1] One specific example is the natural justification in the user simulation framework for the use of variable $k$ when using NDCG@k to aggregate over a set of queries [Karmaker:2020:CIKM].
@@
To promote user simulation for evaluation, as a first step, we should leverage the large number of existing test collections, notably those created at TREC, CLEF, and NCTIR, to turn those into user simulators that are made publicly available to the community. For example, we may use a detailed topic description from a TREC collection as assumed information about a target user to be simulated, based on which we can then build a simulator to attempt to simulate that user’s behaviour. Search log data can also be used as a basis. For example, in the case of product search, we may attempt to simulate a user who is known to have bought a particular product (based on information available in a search log) to simulate the user’s query formulation behaviour. Even though present-day user simulators are far from perfect or accurate, compared with the current evaluation methods used (where unrealistic or extremely naive user behaviour has been assumed), utilizing user simulation for evaluation is still desirable, let alone the opportunity to evaluate interactive systems that we currently cannot easily do without relying on user simulation [Zhang:2017:ICTIR]. [^2] Note that despite the incompleteness of relevance judgments, relative comparison of systems using such imperfect relevance judgments is still meaningful, so in this sense, the community is already accepting this kind of “imperfect evaluation.”
@@
User simulators can be released as toolkits, such as SimIIR [Maxwell:2016:SIGIR, Zerhoudi:2022:CIKM] or UserSimCRS [Afzali:2023:WSDM], or a web-based API service as has been done in living lab-style evaluation [Hopfgartner:2019:bookchapter, Jagerman:2018:JDIQ]. These relevant existing efforts have already demonstrated the feasibility of using more interactive evaluation approaches. To promote the use of user simulation for evaluation, it is also necessary to organize evaluation activities regularly (e.g., by establishing a user simulation track at TREC, CLEF, or NTCIR) to facilitate research and sharing of knowledge.